OpenStudy (anonymous):

Raise this number to its power using the DeMoivre's theorem: (square root3+i)^5 a. 32(cos300degrees+isin300degrees) b. 10(cos150degrees+isin 150degree) c. 243(cos97.5degree+sin 97.5degrees) d. 32(cos150degres+isin150degrees)

OpenStudy (anonymous):

\[\sqrt{3} +i = 2(e^{i \frac{\pi}{6}})\]

OpenStudy (anonymous):

yeah, i have no idea what that is...

OpenStudy (anonymous):

it's the first choice.